# Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence

### v3

@@ Appendix A Contractive Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the contractive loop experiment described in Section 7. The initial artifact $a_{0}$ undergoes 50 iterations of the rewriting transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
The progressive evolution of the text demonstrates the contractive dynamics analyzed in Subsection 7.2: initial rapid refinement (iterations 0–20), followed by stabilization into a dense semantic attractor (iterations 20–50) where only minor stylistic variations occur while the core message remains fixed.
@@ Appendix B Exploratory Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment described in Section 7. The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation, producing the trajectory $\{a_{t}\}_{t=0}^{50}$ shown below.
The trajectory demonstrates the exploratory dynamics analyzed in Subsection 7.3: persistent semantic instability with no convergence, exhibiting large inter-iteration distances and continuous oscillation between opposing philosophical positions without settling into any attractor basin.

### v4

@@ Appendix A Contractive Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the contractive loop experiment. The initial artifact $a_{0}$ undergoes 50 iterations of the rewriting transformation.
The progressive evolution demonstrates contractive dynamics: initial rapid refinement (iterations 0–20), followed by stabilization into a dense semantic attractor (iterations 20–50) where only minor stylistic variations occur while the core message remains fixed.
@@ Appendix B Exploratory Loop: Complete Trajectory
This appendix provides the complete sequence of artifacts generated by the exploratory loop experiment. The initial artifact $a_{0}$ undergoes 50 iterations of a summarize-then-negate transformation.
The trajectory demonstrates exploratory dynamics: persistent semantic instability with no convergence, exhibiting large inter-iteration distances and continuous oscillation between opposing philosophical positions without settling into any attractor basin.